OpenStudy (anonymous):

Just learning Trig. can you help with 2sin^2(2x)=1

OpenStudy (anonymous):

Are you asking with help in solving for x?

OpenStudy (anonymous):

Thats all it said and [0,2 π ]

OpenStudy (fibonaccichick666):

So, did it sound like this: Solve for x contained on \([0,2\pi]\). \(2sin^2(2x)=1\)

OpenStudy (fibonaccichick666):

just to be certain

OpenStudy (anonymous):

it'd be arcsin^2(1/2)

OpenStudy (anonymous):

just said solve over the given interval

OpenStudy (fibonaccichick666):

Alright, so lets start with the bare bones, let's get that sine term by itself on one side, can you do that?

OpenStudy (anonymous):

ok so, divide the 2 on both sides. sin^2(2x)=1 divided by 2 = 1/2

OpenStudy (anonymous):

what do i do with that 2x though?

OpenStudy (fibonaccichick666):

good, now just wait on the 2x

OpenStudy (fibonaccichick666):

now we have \[sin^2(2x)=\frac{1}{2}\]

OpenStudy (fibonaccichick666):

by order of operations, we need to undo the sine function next, how do we 'undo' sine?

OpenStudy (anonymous):

is that when you change it to the numerical equivalent 5?

OpenStudy (fibonaccichick666):

? Not sure what you mean by that, can you elaborate for me please?

OpenStudy (anonymous):

i remember my teacher mentioning that if you put it in the calculator sine is = to .5

OpenStudy (fibonaccichick666):

uhm, no definitely not in this case. and actually, that is a misconception entirely. But I do 100% approve of the effort, so lets think. I'm going to give you something similar to think on. If you have 4x=3, what do you do to solve for x? How about \(x^2y=6\) and you solve for y?

OpenStudy (anonymous):

you would move the 4 to the left side to solve for x

OpenStudy (anonymous):

same with the second one get the y alone

OpenStudy (fibonaccichick666):

Can you show me exactly what you do, like painstakingly step by step?

OpenStudy (anonymous):

divide 4 on the left leaving x and 4 on the right making it 3/4 x=3/4

OpenStudy (fibonaccichick666):

ok so you did this? \[\frac{1}{4} \times 4x=3 \times \frac{1}{4}\] correct?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

In doing so, I am multiplying both sides by the _____.

OpenStudy (fibonaccichick666):

(starts with an i)

OpenStudy (anonymous):

inverse?

OpenStudy (fibonaccichick666):

perfect

OpenStudy (fibonaccichick666):

(that was all I wanted you to take away from those examples)

OpenStudy (fibonaccichick666):

so now, in order to get rid of the sine we need the inverse of sine, do you know what that is?

OpenStudy (anonymous):

no i do not sorry

OpenStudy (fibonaccichick666):

it's ok, that is something that is you either know it, or you do not, so I direct you here to learn about inverse trig functions, just tag me when you get back and we can continue the problem ok? https://www.khanacademy.org/math/trigonometry/unit-circle-trig-func/inverse_trig_functions/v/inverse-trig-functions--arcsin